# Growth rate of liquidity provider’s wealth in G3Ms

### v3

@@ Abstract.
We study how trading fees and continuous-time arbitrage affect the profitability of liquidity providers (LPs) in Geometric Mean Market Makers (G3Ms). We use stochastic reflected diffusion processes to analyze the dynamics of a G3M model [Eva21] under the arbitrage-driven market [MMRZ22a]. Our research focuses on calculating LP wealth and extends the findings of Tassy and White [TW20] related to the constant product market maker (Uniswap v2) to a wider range of G3Ms, including Balancer. This allows us to calculate the long-term expected logarithmic growth of LP wealth, offering new insights into the complex dynamics of AMMs and their implications for LPs in decentralized finance.
@@ 1. Introduction
Decentralized finance (DeFi) has revolutionized financial markets by enabling trustless trading and investment activities through blockchain technology [CJ21, GM23]. At the heart of this revolution are Automated Market Makers (AMMs) [Moh22], which replace traditional order book exchanges with smart contracts that automate price discovery and order execution. This innovation eliminates intermediaries, reduces trading costs, and enhances accessibility for a wider range of participants.
Among the diverse landscape of AMM designs, Geometric Mean Market Makers (G3Ms) have gained significant traction, powering popular DeFi protocols like Uniswap [AZR20] and Balancer [MM19]. G3Ms utilize a weighted geometric mean function to determine asset prices within their liquidity pools. This unique pricing mechanism fosters a predictable relationship between asset reserves and prices, promoting market efficiency and arbitrage resistance.
@@
A large body of literature has extensively examined various aspects of AMM mechanics. [AKC+19, AC20] analyzed the theoretical properties and price-tracking capabilities of constant function markets. [Eva21] extended this analysis to G3Ms with variable weights, while [FMW23, FMW24] investigated impermanent loss and hedging strategies. The concept of "loss-versus-rebalancing" (LVR) was quantified by Milionis et al. [MMRZ22b, MMRZ22a, MMR23], who developed frameworks for measuring systematic LP losses in both standard AMMs and concentrated liquidity markets. [CDM23, CDM24] explored predictable loss in constant function markets and developed optimal liquidity provision strategies under separate fee rate models. [BEK24] analyzed risks and incentives in AMM liquidity provision, proposing novel transaction cost structures. More recently, [NTYY24] examined arbitrage-driven price dynamics in fee-based AMMs, providing insights into the relationship between fee structures and market behavior.
Previous research has extensively examined various aspects of AMM mechanics. Angeris et al. [AKC+19, AC20] analyzed the theoretical properties and price-tracking capabilities of constant function markets. Evans [Eva21] extended this analysis to G3Ms with variable weights, while others like Fukasawa et al. [FMW23, FMW24] investigated impermanent loss and hedging strategies. Milionis et al. [MMRZ22b, MMRZ22a, MMR23], who quantified it as "loss-versus-rebalancing" (LVR) and developed frameworks for measuring systematic LP losses in both standard AMMs and concentrated liquidity markets.
While the existing literature provides a rich foundation for understanding AMMs, a key question remains: How do arbitrage dynamics and trading fees interact to influence the long-term growth of liquidity provider (LP) wealth in G3Ms? This paper aims to address this question by developing a novel framework for analyzing LP wealth evolution in G3Ms, extending beyond the specific case of Uniswap V2 studied in [TW20]. We model the G3M under the influence of continuous arbitrage activity in the presence of a frictionless reference market, drawing inspiration from the framework presented in [MMRZ22b, MMRZ22a, MMR23]. Our analysis leverages the theory of reflected diffusion processes to capture the dynamics of mispricing and its effect on LP returns.
This research makes three key contributions:
@@
    We derive explicit formulas for LP wealth growth rates under various market conditions, encompassing different volatility and drift scenarios.
@@
    We analyze how different fee structures affect LP returns, offering insights into the optimal design of G3Ms.
@@
    We compare the performance of G3Ms to traditional constant rebalanced portfolio strategies, highlighting the potential of G3Ms to serve as an index infrastructure within the DeFi ecosystem.
By employing reflected diffusion processes [LS84], we extend and generalize existing results on LP wealth dynamics, providing a more comprehensive and unified framework. Our findings offer valuable insights for both LPs and AMM designers, contributing to a deeper understanding of DeFi market mechanisms and their implications for investor returns.
@@
The remainder of the paper is organized as follows. Section 2 establishes notation and examines market dynamics with and without transaction costs. Section 3 explores the arbitrageur’s problem and G3M dynamics under both discrete and continuous arbitrage. Section 4 presents our main results, connecting LP wealth growth to parabolic PDEs with Neumann boundary conditions and extending results to cases with independent stochastic volatility and drift.
@@ 2. Constant Weight G3Ms
Geometric Mean Market Makers (G3Ms) are a prominent class of automated market makers (AMMs) that utilize a weighted geometric mean to define the feasible set of trades. This section provides a detailed review of G3M mechanisms, following the work in [Eva21, Moh22, FMW24], for a system with two assets, $X$ and $Y$, and a fixed weight $w\in(0,1)$.
@@ 2.1.1. Trading without Transaction Costs
A key advantage of G3Ms without transaction costs is their path independence [AC20, §2.3]. This means the final state of the pool depends only on the net change in reserves, not the specific sequence of trades that led to it. This property makes these G3Ms resistant to price manipulation strategies.
@@ 2.1.2. With Transaction Costs
    Unlike G3Ms without transaction fees, the presence of transaction costs introduces path dependence [AC20, §2.3]. This means that the order and size of trades influence the final outcome. To illustrate this, consider a trader selling $\Delta_{x}$ of asset $X$ in exchange for $\Delta_{y}$ of asset $Y$. The transaction satisfies
@@
    Comparing these equations reveals that $\Delta_{y}<\Delta^{1}_{y}+\Delta^{2}_{y}<0$, demonstrating that splitting the trade leads to a higher cost (i.e., a smaller amount of asset Y received). This phenomenon is further explored in [AKC+19, Appendix D] and [FMW24, Proposition 1].
@@ 3.1. Arbitrage Bounds and Price Dynamics
To understand these bounds, let’s denote the price of asset $X$ relative to asset $Y$ in the reference market as $S$, and the corresponding price in the G3M pool as $P$. Arbitrageurs aim to maximize their profit by buying asset $X$ where it’s cheaper and selling it where it’s more expensive. This leads to the following optimization problem for an arbitrageur, as shown in [AKC+19]:
@@
where $\gamma$ is the transaction cost parameter of the G3M [AKC+19, §2.1]. This interval, $\gamma S,\gamma^{-1}S]$, defines the no-arbitrage bounds.
@@
Figure 1. Price time series (1-minute intervals) from 14:00:00 to 18:00:00 on September 13, 2023. Dashed lines indicate the upper and lower boundaries of the no-arbitrage region.
@@ 3.2. Mispricing and Arbitrage Dynamics
To formally analyze how arbitrageurs influence the G3M’s behavior, we introduce the concept of mispricing, which quantifies the discrepancy between the G3M pool price and the reference market price. We investigate how this mispricing evolves under both discrete and continuous arbitrage scenarios, building upon the framework in [MMRZ22a, MMR23].
@@ 3.2.1. Discrete Arbitrage Model
Following the discussion of Section 3.1, the arbitrage process can be described by the following recipe, similar to [MMR23, Lemma 2]:
@@
  • Fukasawa et al. [FMW23, FMW24] analyzed arbitrage in G3Ms by focusing on the dynamics of the reserve processes. Here, we adopt a different perspective by directly analyzing the mispricing process, which is compatible with their work through Equation (6). Our approach has the advantage of analyzing the growth rate of liquidity provider wealth, as discussed in Section 4.
@@ 3.2.2. Continuous Arbitrage Model
Following the approach in [FMW24, §3], we decompose the log pool price as $\ln P_{t}=\ln P_{0}+U_{t}-L_{t}$, where $U_{t}$ and $L_{t}$ represent the cumulative changes in log price due to buy and sell arbitrage orders, respectively. These processes naturally satisfy properties (a) to (c). Equations (19) and (20) then follow directly from the characterization of reflected processes in [Har13, Proposition 2.4]. ∎
@@
The mispricing process $Z_{t}$ can be viewed as a stochastic storage system with finite capacity, where $L_{t}$ and $U_{t}$ act as reflecting barriers that keep $Z_{t}$ within the allowed range [Har13, §2.3]. In the special case where the reference market price $S_{t}$ follows a geometric Brownian motion, $Z_{t}$ becomes a reflected Brownian motion [Har13, §6].
We now consider the dynamics of the reserve process $x_{t}$ (or $y_{t}$). Time $t$ is a point of increase (or decrease) for $x_{t}$ if there exists $\delta>0$ such that $x_{t-\delta_{1}}<x_{t+\delta_{2}}$ (or $x_{t-\delta_{1}}>x_{t+\delta_{2}}$) for all $\delta_{1},\delta_{2}\in(0,\delta]$. The reserve process $x_{t}$ increases (or decreases) only when $Z_{t}=a$ if, at every point of increase (or decrease) for $x_{t}$, $Z_{t}=a$.
@@
The liquidity growth term in (22) corresponds to the excess growth rate (see Section 4.1) in Stochastic Portfolio Theory [Fer02, §1.1] as $\gamma\to 1$.
@@ 4.2. Mispricing Process as a Reflected Diffusion
Fix a fee tier $\gamma\in(0,1)$ and set $c=-\ln\gamma$. Suppose that $\ln(S_{0}/P_{0})=z_{0}\in(-c,c)$. According to [Har13, Proposition 2.4], $s_{t}$ uniquely determines a process $Z_{t}$ that satisfies the following properties:
@@
One would anticipate that $L_{t}$ and $U_{t}$ are boundary local times of $Z_{t}$. This is justified by the following lemma. For the basic theory on local times, see [RY99].
@@
The computation of the expectation of (33) simplifies when $\tilde{\mu}$ and $\tilde{\sigma}$ are constant because $Z_{t}$ has Markov property. However, this is not the case when $\tilde{\mu}$ or $\tilde{\sigma}$ depend on $s_{t}$. In order to overcome this technical obstacle, we appeal to the method of mimicking [Kry85, Gyö86] and adapt it to our current setting where local times are involved in the SDE. Define
@@
By Theorem 3.1 and Remark 3.3 of [LS84], the Skorokhod problem
@@
Then, we may use Itô’s lemma as in the proof of Lemma 4.1 to show that $k^{-c}_{t}=\hat{L}_{t}$ and $k^{c}_{t}=\hat{U}_{t}$, and hence (35) holds. Finally, as we know the solution $\hat{Z}_{t}$ to (35) is pathwise unique (because if $\tilde{Z}_{t}$ is another solution, then $(\tilde{Z}_{t},\tilde{U}_{t}-\tilde{L}_{t})$ with $\tilde{L}_{t}$ and $\tilde{U}_{t}$ being boundary local times of $\tilde{Z}_{t}$ is a solution to (38) which is known to be unique), it follows from the standard theory that the solution $\hat{Z}_{t}$ to (35) is weakly unique and hence has strong Markov property; see [Bas98, §I.4–I.5] or [KS91, §5.3–5.4]. ∎

### v4

@@ Abstract.
We study how trading fees and continuous-time arbitrage affect the profitability of liquidity providers (LPs) in Geometric Mean Market Makers (G3Ms). We use stochastic reflected diffusion processes to analyze the dynamics of a G3M model under the arbitrage-driven market [milionis2022automated]. Our research focuses on calculating LP wealth and extends the findings of Tassy and White [tassy2020growth] for the constant product market maker (Uniswap v2) to a broader range of G3Ms, including Balancer. This allows us to calculate the long-term expected logarithmic growth of LP wealth, offering new insights into the complex dynamics of AMMs and their implications for LPs in decentralized finance.
@@ 1. Introduction
Decentralized finance (DeFi) has transformed financial markets by enabling trustless trading and investment through blockchain technology [capponi2021adoption, gobet2023decentralized]. At the heart of this transformations are Automated Market Makers (AMMs) [mohan2022automated], which replace traditional order book exchanges with smart contracts that automate price discovery and trade execution, eliminating intermediaries and broadening market accessibility. Among the diverse landscape of AMM designs, Geometric Mean Market Makers (G3Ms) have gained significant traction, powering leading DeFi protocols such as Uniswap [adams2020uniswap] and Balancer [martinelli2019balancer]. By maintaining a fixed weighted geometric mean of asset reserves, a G3M enforces a constant target allocation across its constituent assets, creating a natural rebalancing dynamic that closely mirrors the discipline of a classical constant-weighted index fund. This structural parallel raises a natural and practically important question: can a G3M, through its fee mechanism, generate long-term wealth growth that matches or even exceeds that of a traditional passive investment strategy?
The goal of this paper is to address this question by developing a tractable analytical framework for studying the long-term growth of liquidity provider (LP) wealth in G3Ms, explicitly accounting for the interplay between continuous-time arbitrage and transaction fees. Our approach is motivated by the observation that, in a G3M, every trade moves the pool price toward the external market price, generating fee revenue for LPs while simultaneously exposing them to adverse selection. Understanding how these two forces interact over long horizons is essential for evaluating G3Ms as investment vehicles and for designing fee structures that maximize LP returns.
We model the G3M under the influence of continuous arbitrage activity against a frictionless external reference market, adapting the arbitrage-driven price dynamics of [najnudel2024arbitrage] and drawing on the loss-versus-rebalancing framework of [milionis2022quantifying, milionis2022automated, milionis2023automated]. The key analytical tool is the theory of reflected diffusion processes [lions1984stochastic]: the log-ratio of the external market price to the G3M pool price, which we call the mispricing process, evolves as a diffusion confined to the no-arbitrage interval $\ln\gamma,-\ln\gamma]$ by reflecting boundary local times that encode the fee-driven arbitrage activity. The long-term growth rate of LP wealth can then be expressed in terms of the ergodic properties of this reflected diffusion, yielding explicit formulas that connect fee structures, market volatility, and portfolio weights.
The scope of this paper is not to provide a fully realistic microstructural model of AMM trading, but to establish a clean baseline that isolates the arbitrage channel and demonstrates the potential of G3Ms to serve as on-chain index fund infrastructure. In particular, our results show that, from a portfolio construction perspective, short-term arbitrage losses can be transformed into long-run volatility harvesting through the fee mechanism, with the LP wealth growth rate outperforming the excess growth rate of a constant-rebalanced portfolio in Stochastic Portfolio Theory (SPT) [fernholz2002stochastic, karatzas2009stochastic] — which itself assumes frictionless trading — when the fee tier is chosen appropriately.
To this end, we adopt three simplifying assumptions, each of which we now motivate.
Assumption A: No Noise Traders.
We assume all trades are executed by arbitrageurs, with no noise traders present. This is a deliberate worst-case baseline: in practice, retail order flow is typically routed through aggregation solvers such as 1inch or CowSwap, whose highly nonlinear routing logic is difficult to incorporate into a closed-form model (see, e.g., [brady2024amm]). Under this assumption, LPs face only adversarial order flow and are in the least favorable scenario for fee collection. Nevertheless, our results demonstrate that even in this worst-case setting, the LP wealth growth rate can outpace that of a constant-rebalanced portfolio under SPT, provided the fee tier is chosen appropriately. The presence of uninformed order flow would, in practice, only strengthen this conclusion.
Assumption B: Frictionless External Reference Market.
We assume the existence of an external reference price with infinite liquidity and no transaction costs, representing an aggregated oracle across all available liquidity sources — DEXs, CEXs, and routing solvers — rather than the price on any single venue. The special case of a single CEX with proportional transaction costs is studied in [campbell2025optimal], where the arbitrage structure takes a similar form to ours but with a modified no-arbitrage boundary; in that setting, our growth rate computation carries over verbatim. Incorporating fixed transaction costs, however, would break the scale homogeneity of the mispricing dynamics, substantially increasing analytical complexity. We therefore leave a systematic treatment of inter-venue transaction costs as an important direction for future work.
Assumption C: Continuous Arbitrage.
We assume arbitrageurs continuously monitor and act on arbitrage opportunities instantaneously. While idealized, this assumption is increasingly justified by modern blockchain infrastructure: block times on networks such as Arbitrum ($\approx 250$ms), Solana ($\approx 400$ms), and MegaETH ($\approx 10$ms) are short enough that the continuous-time limit provides a reasonable approximation of high-frequency arbitrage activity.
1.1. Literature Review
AMMs.
@@ 1.1. Literature Review
A large body of literature has examined various aspects of AMM mechanics. [angeris2019analysis, angeris2020improved] analyzed the theoretical properties and price-tracking capabilities of constant function markets. [evans2021liquidity] extended this analysis to G3Ms with variable weights, while [fukasawa2023weighted, fukasawa2023modelfree] investigated impermanent loss and hedging strategies. More recently, [najnudel2024arbitrage] examined arbitrage-driven price dynamics in fee-based AMMs, providing insights into the relationship between fee structures and market behavior.
LP Risk.
A central concern in the AMM literature is the quantification and management of LP losses. [milionis2022quantifying, milionis2022automated, milionis2023automated] introduced the concept of “loss-versus-rebalancing” (LVR) and developed frameworks for measuring systematic LP losses in both standard AMMs and concentrated liquidity markets. [cartea2023predictable, cartea2024decentralized] explored predictable loss in constant function markets and developed optimal liquidity provision strategies under separate fee rate models. [bronnimann2024automated] further analyzed risks and incentives in AMM liquidity provision, proposing novel transaction cost structures. [risk2026pricing] develops a pricing and hedging framework for CFMMs, characterizing impermanent loss as a weighted strip of vanilla options via the Carr–Madan spanning formula and validating the resulting implied volatility structure against Uniswap v3 and Deribit option data. Beyond loss quantification, [drissi2025equilibrium] studied equilibrium liquidity provision in decentralized markets, demonstrating how LPs can offset inventory and adverse selection risks through hedging strategies, and highlighting the importance of fee income in the context of LP risk-return trade-offs.
Blockchain Market Microstructure.
A growing literature examines microstructural features specific to blockchain trading environments that are absent from classical market models. Factors such as gas costs, block discreteness, and market fragmentation can materially affect LP returns and arbitrage dynamics. [aoyagi2025coexisting] examines the interaction between AMMs and centralized limit order books, demonstrating how their coexistence influences liquidity provision and price discovery. [capponi2025viability] develops a model to evaluate the viability of blockchain markets as the sole venue for price formation, with particular attention to the effects of block times on market efficiency. Closely related to our analysis of optimal fee tiers, [campbell2025optimal] studies LP profitability in a dynamic model where an AMM operates alongside a CEX with optimal order routing and arbitrage, finding that the optimal fee is stable under normal market conditions but should be raised during periods of high volatility — a finding that complements our characterization of the fee tier that maximizes long-term LP wealth growth.
1.2. Contribution
This paper extends the analysis of [tassy2020growth], which studied LP wealth growth in the constant product market maker (Uniswap V2), to a broad class of G3Ms, including Balancer. Within the framework described above, the paper makes three key contributions:
@@ 1.2. Contribution
    We derive explicit formulas for the long-term expected logarithmic growth rate of LP wealth under various market conditions, including time-homogeneous, time-inhomogeneous, and stochastic volatility settings, and across different volatility and drift scenarios.
@@
    We analyze how different fee structures affect LP returns, characterizing the optimal fee tier that maximizes LP wealth growth and revealing non-monotone phase transitions as a function of pool weights and market drift.
@@
    We compare the long-term performance of G3Ms to constant-rebalanced portfolio strategies in the sense of SPT, demonstrating that G3Ms can serve as a competitive on-chain index fund infrastructure capable of outperforming classical passive investment benchmarks.
@@
The remainder of the paper is organized as follows. Section 2 establishes notation and examines G3M market dynamics with and without transaction costs. Section 3 studies the arbitrageur’s problem and G3M dynamics under both discrete and continuous arbitrage. Section 4 presents our main results, connecting LP wealth growth to parabolic PDEs with Neumann boundary conditions and extending the analysis to cases with independent stochastic volatility and drift.
@@ 2. Constant Weight G3Ms
Geometric Mean Market Makers (G3Ms) are a prominent class of automated market makers (AMMs) that utilize a weighted geometric mean to define the feasible set of trades. This section provides a detailed review of G3M mechanisms, following the work in [evans2021liquidity, mohan2022automated, fukasawa2023modelfree], for a system with two assets, $X$ and $Y$, and a fixed weight $w\in(0,1)$.
@@ 2.1.1. Trading without Transaction Costs
A key advantage of G3Ms without transaction costs is their path independence [angeris2020improved, §2.3]. This means the final state of the pool depends only on the net change in reserves, not the specific sequence of trades that led to it. This property makes these G3Ms resistant to price manipulation strategies.
@@ 2.1.2. With Transaction Costs
    Unlike G3Ms without transaction fees, the presence of transaction costs introduces path dependence [angeris2020improved, §2.3]. This means that the order and size of trades influence the final outcome. To illustrate this, consider a trader selling $\Delta_{x}$ of asset $X$ in exchange for $\Delta_{y}$ of asset $Y$. The transaction satisfies
@@
    Comparing these equations reveals that $\Delta_{y}<\Delta^{1}_{y}+\Delta^{2}_{y}<0$, demonstrating that splitting the trade leads to a higher cost (i.e., a smaller amount of asset Y received). This phenomenon is further explored in [angeris2019analysis, Appendix D] and [fukasawa2023modelfree, Proposition 1].
@@ 3.1. Arbitrage Bounds and Price Dynamics
To understand these bounds, let’s denote the price of asset $X$ relative to asset $Y$ in the reference market as $S$, and the corresponding price in the G3M pool as $P$. Arbitrageurs aim to maximize their profit by buying asset $X$ where it’s cheaper and selling it where it’s more expensive. This leads to the following optimization problem for an arbitrageur, as shown in [angeris2019analysis]:
@@
where $\gamma$ is the transaction cost parameter of the G3M [angeris2019analysis, §2.1]. This interval, $\gamma S,\gamma^{-1}S]$, defines the no-arbitrage bounds.
@@
Figure 1. Price time series (1-minute intervals) from 14:00:$0$ to 18:00:$0$ on September 13, 2023. Dashed lines indicate the upper and lower boundaries of the no-arbitrage region.
@@ 3.2. Mispricing and Arbitrage Dynamics
To formally analyze how arbitrageurs influence the G3M’s behavior, we introduce the concept of mispricing, which quantifies the discrepancy between the G3M pool price and the reference market price. We investigate how this mispricing evolves under both discrete and continuous arbitrage scenarios, building upon the framework in [milionis2022automated, milionis2023automated].
@@ 3.2.1. Discrete Arbitrage Model
Following the discussion of Section 3.1, the arbitrage process can be described by the following recipe, similar to [milionis2023automated, Lemma 2]:
@@
  • Fukasawa et al. [fukasawa2023weighted, fukasawa2023modelfree] analyzed arbitrage in G3Ms by focusing on the dynamics of the reserve processes. Here, we adopt a different perspective by directly analyzing the mispricing process, which is compatible with their work through Equation (6). Our approach has the advantage of analyzing the growth rate of liquidity provider wealth, as discussed in Section 4.
@@ 3.2.2. Continuous Arbitrage Model
Following the approach in [fukasawa2023modelfree, §3], we decompose the log pool price as $\ln P_{t}=\ln P_{0}+U_{t}-L_{t}$, where $U_{t}$ and $L_{t}$ represent the cumulative changes in log price due to buy and sell arbitrage orders, respectively. These processes naturally satisfy properties (a) to (c). Equations (19) and (20) then follow directly from the characterization of reflected processes in [harrison2013brownian, Proposition 2.4]. ∎
@@
The mispricing process $Z_{t}$ can be viewed as a stochastic storage system with finite capacity, where $L_{t}$ and $U_{t}$ act as reflecting barriers that keep $Z_{t}$ within the allowed range [harrison2013brownian, §2.3]. In the special case where the reference market price $S_{t}$ follows a geometric Brownian motion, $Z_{t}$ becomes a reflected Brownian motion [harrison2013brownian, §6].
Definition 3.9.
We say time $t$ is a point of increase (resp. point of decrease) for $x_{t}$ if there exists $\delta>0$ such that $x_{t-\delta_{1}}<x_{t+\delta_{2}}$ (resp. $x_{t-\delta_{1}}>x_{t+\delta_{2}}$) for all $\delta_{1},\delta_{2}\in(0,\delta]$. We say that $x_{t}$ increases (resp. decreases) only when $Z_{t}=a$ if, at every point of increase (resp. point of decrease) of $x_{t}$, we have $Z_{t}=a$.
@@
The liquidity growth term in (22) corresponds to the excess growth rate (see Section 4.1) in Stochastic Portfolio Theory [fernholz2002stochastic, §1.1] as $\gamma\to 1$.
@@ 4.2. Mispricing Process as a Reflected Diffusion
Fix a fee tier $\gamma\in(0,1)$ and set $c=-\ln\gamma$. Suppose that $\ln(S_{0}/P_{0})=z_{0}\in(-c,c)$. According to [harrison2013brownian, Proposition 2.4], $s_{t}$ uniquely determines a process $Z_{t}$ that satisfies the following properties:
@@
One would anticipate that $L_{t}$ and $U_{t}$ are boundary local times of $Z_{t}$. This is justified by the following lemma. For the basic theory on local times, see [revuz1999continuous].
@@
The computation of the expectation of (33) simplifies when $\tilde{\mu}$ and $\tilde{\sigma}$ are constant because $Z_{t}$ has Markov property. However, this is not the case when $\tilde{\mu}$ or $\tilde{\sigma}$ depend on $s_{t}$. In order to overcome this technical obstacle, we appeal to the method of mimicking [krylov1985once, gyongy1986mimicking] and adapt it to our current setting where local times are involved in the SDE. Define
@@
By Theorem 3.1 and Remark 3.3 of [lions1984stochastic], the Skorokhod problem
@@
Then, we may use Itô’s lemma as in the proof of Lemma 4.1 to show that $k^{-c}_{t}=\hat{L}_{t}$ and $k^{c}_{t}=\hat{U}_{t}$, and hence (35) holds. Finally, as we know the solution $\hat{Z}_{t}$ to (35) is pathwise unique (because if $\tilde{Z}_{t}$ is another solution, then $(\tilde{Z}_{t},\tilde{U}_{t}-\tilde{L}_{t})$ with $\tilde{L}_{t}$ and $\tilde{U}_{t}$ being boundary local times of $\tilde{Z}_{t}$ is a solution to (38) which is known to be unique), it follows from the standard theory that the solution $\hat{Z}_{t}$ to (35) is weakly unique and hence has strong Markov property; see [bass1998diffusions, §I.4–I.5] or [karatzas1991brownian, §5.3–5.4]. ∎